The substance (CH_3CH_2)_2NH is considered Which
Answer #1

15)

this is diethyl amine; its a pretty common WEAK BASE since it won't ionize completely

16)

OH- and O2-

since O2- is not actually in solution

OH- may only gain H+, therefore its conjugate will be H2O not O-2

17)

The soluition is basic since

NaF --> Na+ and F-

F- + H2O <-> HF + OH-

18)

Buffer is a weak base/acid + conjugate acid/base respectively so:

none of these is the case

19)

for a pH = 4.5 choose a pKa = pH or nearest

or pOH = 14-4.5 = 9.5 or pKb = 9.5 or nearest

choose B since Ka = 1.8*10^-5 then pK a= -log(1.8*10^-5) = 4.7447

20)

Adding KOH will increase OH- ions, therefore the equilibrium will shift toward the reactants

the solubility will DECREASE since more solid product is favoured